An Indian multinational corporation has a subsidiary in the United Kingdom. To meet its investment needs that subsidiary firm wants to raise GBP 10 million by issuing five-year bonds in the British market at a coupon rate of 10percent per annum, payable annually. The current spot rate is GBP/INR 82. The INR is expected to depreciate at 1 percent each year for the next five years what is the effective cost of debt to the Indian MNC? Determine the Net Cash Flows for five years is as follows. 8marks


Year 1 2 3 4 5


57.98 62.75 63.38 64.02 926.93



1
Expert's answer
2022-03-07T18:08:33-0500

"total debt=10\times82+10\times82\times0.1\times5=1230"


"the effective cost of debt=\frac{1230-57.98-62.75-63.3- 64.02-926.93}{1230}\times100=4.49"